# Charge transport and mode transition in dual-energy electron beam diodes

### v2

@@ Abstract
Introduction—Electron beams Tsimring (2007); Go (2022); Sharma and Hara (2025), as directed electron streams characterized by their current density Li et al. (2025); Campanell et al. (2012); Campanell and Umansky (2016), energy spectrum Wenz et al. (2019); Furman and Pivi (2002); Chen and Fu (2025a) and spatiotemporal evolutions Heri and Zhang (2025); Ang and Zhang (2007), are indispensable drivers for various applications Zhang et al. (2017); Pagán Muñoz et al. (2024); Dubrovski et al. (2024); Chen and Fu (2025b), such as virtual cathode oscillators Jiang (2024); Mumtaz et al. (2024), thermionic energy converters Sitek et al. (2021); Abdul Khalid et al. (2016); Lin et al. (2024), compact terahertz sources Pedersen et al. (2010); Zhang et al. (2025); Kumar et al. (2025), and ultrafast switches Samizadeh Nikoo et al. (2020). In electron beam driven diodes, an injected intense electron current, accompanied by significant space charge effects Birdsall and Bridges (1966); Campanell et al. (2025), can limit the transmitted current to a maximum, which is known as the space-charge-limited (SCL) current $J_{\text{SCL}}$ Child (1911); Jaffé (1944). The charge transport in diodes with mono-energetic electron beams has been well understood, and the maximum transmitted current obeys the prediction by the SCL current $J_{\text{SCL}}$ Zhu et al. (2013, 2023); Garner et al. (2025). Previously, first-principle particle-in-cell (PIC) simulations showed that in mono-energetic electron beam driven diodes, the transmitted current $J_{\text{tran}}$ is as a function of the injected current $J_{\text{inj}}$, i.e., $J_{\text{tran}}=J_{\text{inj}}$ for $J_{\text{inj}}\leq J_{\text{SCL}}$ whereas $J_{\text{tran}}$ becomes saturated (or oscillating with nonzero emission velocity) for $J_{\text{inj}}>J_{\text{SCL}}$ Lafleur (2020, 2022). If $J_{\text{inj}}>J_{\text{SCL}}$, a virtual cathode (adjacent to the cathode) can form with an oscillating potential barrier, which periodically reflects less energetic electrons released from the cathode, thereby inducing space-charge oscillations Bridges and Birdsall (1963); Lin and Fu (2025); Zhu and Kong (2023). In the oscillation regime, existing analytical solutions can predict the spatial profiles of diode physical quantities (e.g., electric potential) Lafleur (2020). The scaling law for the space charge oscillation frequency has been theoretically derived and validated via PIC simulations Lin et al. (2025). However, the SCL current fails when electrons exhibit distributed energies (e.g., the Maxwell distribution for electron thermal motion) or multi-energy electron beams can form, since the space charge oscillations can be damped, resulting in a new steady state Lafleur (2020). Although the SCL current and oscillatory dynamics in mono-energy electron beam diodes have been well studied, electron beams with discrete energy components (e.g., dual-energy electron beams) exhibit emergent complexities, remaining to be understood for an uncharted theoretical framework.
@@
In this Letter, we report five complete charge transport modes and their transitions in diodes driven by precisely controlled dual-energy electron beams. For the first time, we, via PIC simulations, demonstrate the control of five distinct operation modes by independently modulating high- and low-energy electron beams, and identify that the velocity ratio dictates mode transitions, whereas the injected current ratio determines transmitted current scaling characteristics.
@@
Table 1: Operation modes in the dual-energy electron beam driven diodes. The low $(\text{e}_{1})$ and high-energy $(\text{e}_{2})$ electrons have three possible modes: transported (T), oscillation (O), and reflected (R) modes.
@@
By separately tuning the control parameters for dual-energy electron beams, the transmitted current scaling and charge transport mode transition are determined, as shown in Fig. 3. The ratio of initial velocity for two electron beams, $\beta_{1}/\beta_{2}$, plays a significant role in determining the charge transport mode transition. For electron beams with a large difference in velocity (energy) (e.g., $\beta_{1}/\beta_{2}=0.1$), the transmitted current characteristics are illustrated in Fig. 3(a). When the injected current $J_{0}$ is sufficiently low (below the SCL current of the low-energy electrons $J_{\text{SCL,1}}$), both low- and high-energy electrons can be fully transmitted [mode M1]. When the injected current for low-energy electrons $J_{1}$ exceeds their corresponding SCL current $J_{\text{SCL,1}}$, the transport of low-energy electrons $\text{e}_{1}$ is suppressed while the high-energy electrons $\text{e}_{2}$ can still be transported completely [mode M2]. Notably, in this regime, the transmitted current for low-energy electrons $J_{\text{tran,1}}$ falls below $J_{1}$ and the reflected current for low-energy electrons $J_{\text{ref,1}}$ increases substantially, maintaining the current conservation $J_{1}=J_{\text{tran,1}}+J_{\text{ref,1}}$. As the injected current further increases, low-energy electrons can be fully reflected with $J_{\text{ref,1}}=J_{\text{inj,1}}$, while high-energy electrons remain fully transported [mode M3]. Once the injected current for high-energy electrons $J_{2}$ is larger than their corresponding SCL current $J_{\text{SCL,2}}$, high-energy electrons $\text{e}_{2}$ are partially reflected, accompanied with the entire reflection of $\text{e}_{1}$ [mode M4]. Consequently, while increasing the total injected current $J_{0}$, the charge transport modes undergo a well-defined pathway: $\text{M}1\rightarrow\text{M}2\rightarrow\text{M}3\rightarrow\text{M}4$, which can be abstracted to discrete eigen-states defined by combinatorial permutations of the transport modes for both energy components [as illustrated in Fig. 3(e)].
@@
Here, in this work five operation modes $(N_{\text{mode}}=5)$ are determined for $n=2$. If one considers identical injected current for nn–component electron beams, $J_{1}=J_{2}=...=J_{n}=J_{0}/n$, with substantial energy differences. The transmitted current $J_{\text{tran}}$ versus the total injected current $J_{0}$ exhibits a stepwise growth [see Fig. 4], and the transmitted current for $\text{e}_{i}$ reaches its maximum $J_{\text{tran},i}=J_{0}/n=J_{\text{SCL},i}$ when $J_{i}=J_{\text{SCL},i}$. If $J_{i}>J_{\text{SCL},i}$ for $\text{e}_{i}$, the transmitted component progressively decreases to zero, and the mode transition $\text{T}\rightarrow\text{O}\rightarrow\text{R}$ occurs. Generally, with an increasing $J_{0}$, the intensified space-charge effects lead to complete reflections of electron beams $\text{e}_{i}$ in the energy sequence, showing stepwise growth in transmitted current. The total transmitted current can be described as
@@
Conclusion—In conclusion, we explicitly determined five complete modes for charge transport in dual-energy electron beam driven diodes, which are modulated by controlling the two electron beam parameters. The initial velocity ratio and the injected current ratio are identified as governing parameters; the former determines the charge transport modes and their transitions, while the latter modulates the characteristics of the transmitted current. The discovery provides a fundamental advance in multi-energy beam physics, enabling precise tuning strategies for electron beam driven diodes, especially for applications in modern vacuum electronic devices.
Acknowledgments—The authors acknowledge financial support from the Beijing Natural Science Foundation of China (No. 3244040) and the Organized Research Support Program (No. YK20240103) from the Department of Electrical Engineering at Tsinghua University.
Data availability—The data that support the findings of this study are available from the contact author upon reasonable request.

### v3

@@ Abstract
Introduction—Electron beams Tsimring (2007); Go (2022); Sharma and Hara (2025), as directed electron streams characterized by their current density Li et al. (2025); Campanell et al. (2012); Campanell and Umansky (2016), energy spectrum Wenz et al. (2019); Furman and Pivi (2002); Chen and Fu (2025a) and spatiotemporal evolutions Heri and Zhang (2025); Ang and Zhang (2007), are indispensable drivers for various applications Zhang et al. (2017); Pagán Muñoz et al. (2024); Dubrovski et al. (2024); Chen and Fu (2025b), such as Pierce diodes Ender et al. (2000), virtual cathode oscillators Jiang (2024); Mumtaz et al. (2024), thermionic energy converters Sitek et al. (2021); Abdul Khalid et al. (2016); Lin et al. (2024), compact terahertz sources Chen et al. (2026); Alexandersson et al. (2026); Pedersen et al. (2010); Zhang et al. (2025); Kumar et al. (2025), and ultrafast switches Samizadeh Nikoo et al. (2020). In electron beam driven diodes, an injected intense electron current, accompanied by significant space charge effects Birdsall and Bridges (1966); Campanell et al. (2025), can limit the transmitted current to a maximum, which is known as the space-charge-limited (SCL) current $J_{\text{SCL}}$ Child (1911); Jaffé (1944). The charge transport in diodes with mono-energetic electron beams has been well understood, and the maximum transmitted current obeys the prediction by the SCL current $J_{\text{SCL}}$ Zhu et al. (2013, 2023); Garner et al. (2025). Previously, first-principle particle-in-cell (PIC) simulations showed that in mono-energetic electron beam driven diodes, the transmitted current $J_{\text{tran}}$ is as a function of the injected current $J_{\text{inj}}$, i.e., $J_{\text{tran}}=J_{\text{inj}}$ for $J_{\text{inj}}\leq J_{\text{SCL}}$, whereas $J_{\text{tran}}$ becomes saturated (or oscillating with nonzero emission velocity) for $J_{\text{inj}}>J_{\text{SCL}}$ Lafleur (2020, 2022). If $J_{\text{inj}}>J_{\text{SCL}}$, a virtual cathode (adjacent to the cathode) can form with an oscillating potential barrier, which periodically reflects less energetic electrons released from the cathode, thereby inducing space-charge oscillations Bridges and Birdsall (1963); Lin and Fu (2025); Zhu and Kong (2023). In the oscillation regime, existing analytical solutions can predict the spatial profiles of diode physical quantities (e.g., electric potential) Lafleur (2020). The scaling law for the space charge oscillation frequency has been theoretically derived and validated via PIC simulations Lin et al. (2025). However, the SCL current fails when electrons have distributed energies (e.g., the Maxwell distribution for electron thermal motion) or when multi-energy electron beams form, since space-charge oscillations can be altered, leading to different steady states Lafleur (2020). Although well-studied in mono-energy beams, the SCL current and oscillatory dynamics in diodes with discrete energy components present emergent complexities, and a complete theoretical framework remains elusive.
@@
In this Letter, we report five complete charge transport modes and their transitions in diodes driven by precisely controlled dual-energy electron beams. For the first time, we, via PIC simulations, demonstrate the control of five distinct operation modes by independently modulating high- and low-energy electron beams, and identify that the velocity ratio dictates mode transitions, whereas the injected current ratio determines transmitted current scaling characteristics. The results provide key insights into multi-electron beam-driven diodes, which are essential for promoting modern technologies in vacuum electronics.
@@
Table 1: Operation modes in the dual-energy electron beam driven diodes. 11footnotetext: Default journal substyle.
@@
By separately tuning the control parameters for dual-energy electron beams, the transmitted current scaling and charge transport mode transition are determined, as shown in Fig. 3. The ratio of initial velocity for two electron beams, $\beta_{1}/\beta_{2}$, plays a significant role in determining the charge transport mode transition. For electron beams with a large difference in velocity (energy) (e.g., $\beta_{1}/\beta_{2}=0.1$), the transmitted current characteristics are illustrated in Fig. 3(a). When the injected current $J_{0}$ is sufficiently low (below the SCL current of the low-energy electrons $J_{\text{SCL,1}}$), both low- and high-energy electrons can be fully transmitted [mode M1]. When the injected current for low-energy electrons $J_{1}$ exceeds their corresponding SCL current $J_{\text{SCL,1}}$, the transport of low-energy electrons $\text{e}_{1}$ is suppressed while the high-energy electrons $\text{e}_{2}$ can still be transported completely [mode M2]. Notably, in this regime, the transmitted current for low-energy electrons $J_{\text{tran,1}}$ falls below $J_{1}$ and the reflected current for low-energy electrons $J_{\text{ref,1}}$ increases substantially, maintaining the current conservation $J_{1}=J_{\text{tran,1}}+J_{\text{ref,1}}$. As the injected current further increases, low-energy electrons can be fully reflected with $J_{\text{ref,1}}=J_{\text{inj,1}}$, while high-energy electrons remain fully transported [mode M3]. Once the injected current for high-energy electrons $J_{2}$ is larger than their corresponding SCL current $J_{\text{SCL,2}}$, high-energy electrons $\text{e}_{2}$ are partially reflected, accompanied with the entire reflection of $\text{e}_{1}$ [mode M4]. Consequently, while increasing the total injected current $J_{0}$, the charge transport modes undergo a well-defined pathway: $\text{M}1\rightarrow\text{M}2\rightarrow\text{M}3\rightarrow\text{M}4$, which can be abstracted to discrete eigen-states defined by combinatorial permutations of the transport modes for both energy components [see Fig. 3(e)].
@@
Here, five operation modes $(N_{\text{mode}}=5)$ are determined for $n=2$. If one considers identical injected current for nn–component electron beams, $J_{1}=J_{2}=...=J_{n}=J_{0}/n$, with substantial energy differences. The transmitted current $J_{\text{tran}}$ versus the total injected current $J_{0}$ exhibits a stepwise growth [see Fig. 4], and the transmitted current for $\text{e}_{i}$ reaches its maximum $J_{\text{tran},i}=J_{0}/n=J_{\text{SCL},i}$ when $J_{i}=J_{\text{SCL},i}$. If $J_{i}>J_{\text{SCL},i}$ for $\text{e}_{i}$, the transmitted component progressively decreases to zero, and the mode transition $\text{T}\rightarrow\text{O}\rightarrow\text{R}$ occurs. Generally, with an increasing $J_{0}$, the intensified space-charge effects lead to complete reflections of electron beams $\text{e}_{i}$ in the energy sequence, showing stepwise growth in transmitted current. The total transmitted current can be described as
@@
Conclusion—In conclusion, we explicitly determined five complete modes for charge transport in dual-energy electron beam driven diodes, which are modulated by controlling the two electron beam parameters. The initial velocity ratio and the injected current ratio are identified as governing parameters; the former determines the charge transport modes and their transitions, while the latter modulates the characteristics of the transmitted current. The discovery provides a fundamental advance in multi-energy beam physics, enabling precise tuning strategies for electron beam-driven diodes, especially for applications in modern vacuum electronic devices.
Acknowledgments—The authors acknowledge the financial support from the Fundamental and Interdisciplinary Disciplines Breakthrough Plan of the Ministry of Education of China (Grant No. JYB2025XDXM312).
Data availability—The data that support the findings of this study are available from the authors upon reasonable request.